# Robust and Dexterous Dual-arm Tele-Cooperation using Adaptable Impedance Control

### 06 Feb 2024 v3

@@ I Introduction
Either admittance or impedance controllers have been integrated into optimised controlled architectures and/or coupled with learning algorithms to implement adaptable controllers [10, 14, 1, 5, 18]. Although these methods successfully implemented proof of concept applications, they are susceptible to lack of accurate models, misrepresentation of the interaction conditions, numerical instability, sensor noise and information delay [8, 19]. Furthermore, they are task-dependent, computationally expensive, and require extensive task-specific parameters tuning, limiting their flexibility [17, 16]. In contrast, animals are capable of dexterous robust interaction in challenging environment despite their bio-signals are noisy and transferred with substantial delays to withstand adverse conditions. Furthermore, they are also capable of learning, adapt, and transfer motor skills, which have fuelled the interest of both the robotics and neuro-scientific communities for the last few decades [10, 20, 2, 3, 4].
@@ III-C Multi-Arm Implementation
However, our earlier experiments validated the superimposition either for redundant manipulator [21, 23] or coordination of a swarm of drones [22]. The proposed method extends the experimental validation of the FIC superimposition properties to multiple redundant manipulators, allowing two arms to complete two bi-manual tasks during unknown interaction. The validation of these properties substantially simplifies the algorithmic complexity of the robot interaction controllers deployed in teleoperation and manipulation. Currently, multi-arm manipulation requires the simultaneous observation of the state of both arms through the grasp matrix, which couples the state of the two systems and increasing the algorithmic complexity [25, 26]. In contrast, the proposed architecture allows solving the same problem just considering the simplified dynamics generated by the FIC at the end-effector.
@@ V Results
Enables coordination of multiple arms using synchronous via-points. |
@@
Stability often requires intricate design and tuning. |
@@
Robust in highly variable environments without re-tuning. |
Might require re-tuning in changing environments. |
@@
Might require additional energy tanks or mechanisms for stability. |
Validation Scenarios |
Validated in diverse scenarios including drilling and human interaction. |
Typically validated in specific, isolated scenarios. |
@@
The error recorded from two cooperation experiments for drilling are the experiments in Fig. 4. They indicate that we can expect sub-millimetre accuracy when the user is holding the drill. The experiment 5.1 proves that the two arm movement can be synchronised issuing coordinated via-points without relying on trajectory planning. The results described in Fig. 5 indicate that it is possible to pick up the work-piece, carry it to the user, providing sufficient support to execute an accurate drilling, and retain stability and safety of interaction even in the event of a sudden loss of contact. The experiment 5.2 confirms the robot can constraint the dremel inclination to the predetermined range of in each direction when drilling on a curve surface, as shown in Fig. 3.
The bi-manual manipulation task trajectories (Exp. 6) are shown in Fig. 6. The arms start in the first via point and move to third to pick-up the box. The empty box is brought in to the 6 via point where the human loads the box with about of material. Regardless of the added mass the robots retain a sufficient position accuracy that enables them to carry the box back to the 3 via point, which is bounded by the FIC force saturation to a value that is . It is worth noting that the FIC is generating the maximum effort (i.e., forces and torques) selected for a task, and any position error greater than implies that the current load exceed the controller/robot capabilities. Thus, it will either require a change in the selected maximum effort in the controller, or a change of robot when the load exceed its mechanical capabilities. Experimental video is accessible through here.
@@ VI Discussion
The tasks show that the method can exploit its nonlinear impedance to generate higher interaction forces and accurate tracking without need of accurate modelling of the interaction dynamics. This characteristics renders this method extremely robust because it does not requires the knowledge of the environmental dynamics to stabilise its behaviour. Therefore, it presents all of the benefits that constant impedance controllers have on variable impedance and admittance controllers, but it does not suffer their low accuracy [17, 10, 1, 27]. Such accomplishment is made possible by the tunable non-linear impedance profile that enables to select the required task precision, the maximum effort and/or constant stiffness that the controller produce. Moreover, the selection of these characteristics also upper-bounds the system maximum power and maximum kinetic energy as described in [24, 22].
The nonlinear impedance allows to define low impedance funnels that can be used to guide the task performed by the user, as it has been shown in Exp. 5.2 when the user was drilling a curved work-piece. This capability might find application in training on new operators, guiding them on specific action patters, or respecting constraints that are otherwise difficult to respect (e.g. drilling with a constant angle on a curve surface). Furthermore, it is worth noting that the interaction with the robot during the task felt very natural and all the tasks required were successful at the first attempt, and the motor synchronisation was obtained only via the mechanical interaction between the robot and the users. Finally, to better understand the unique contributions and advantages of the TeleCoop-FIC, we present a direct comparison with traditional variable impedance controllers that can be seen in Table I. Possible applications for the TeleCoop-FIC can be considered manufacturing, space robotics, medical robotics, rehabilitation and training.
@@ VII Conclusion
In Conclusion, the main observed limitation is that the task accuracy and precision should be in the order of millimetre based on our set-up, but it will vary based on the robots used. Especially considering that the tuning of the controller is straight forward and can be performed by non-technical operators, simply selecting the desired task precision (), maximum effort () with respect to the physical limitation of the robot, and constant stiffness (). However, an identification of the maximum parameters is required to limit the parameter range to be within the mechanical capabilities of the robot. An identification procedure for this type of controller has been already developed, tested on multiple robots and is detailed in [17]. Future research should prioritize refining the controller’s tuning identification procedure, making it adaptable across a diverse range of robot types. Emphasis should also be placed on enhancing the system’s accuracy and precision, addressing the current millimeter-order precision limitations. Additionally, the system’s vast potential in sectors like manufacturing, space robotics, medical robotics, rehabilitation, and training beckons dedicated exploration. Tailoring and optimizing the system for these specific applications can usher in a new era of specialized robotic solutions.

### 29 Feb 2024 v6

@@ I Introduction
Either admittance or impedance controllers have been integrated into optimised controlled architectures and/or coupled with learning algorithms to implement adaptable controllers [10, 14, 1, 5, 18]. Although these methods successfully implemented proof of concept applications, they are susceptible to lack of accurate models, misrepresentation of the interaction conditions, numerical instability, sensor noise and information delay [8, 19]. Furthermore, they are task-dependent, computationally expensive, and require extensive task-specific parameters tuning, limiting their flexibility [17, 16]. In contrast, animals are capable of dexterous robust interaction in challenging environment despite their bio-signals are noisy and transferred with substantial delays to withstand adverse conditions. Furthermore, they are also capable of learning, adapt, and transfer motor skills, which have fuelled the interest of both the robotics and neuro-scientific communities for the last few decades [21, 2, 3, 4].
@@ III-C Multi-Arm Implementation
However, our earlier experiments validated the superimposition either for redundant manipulator [21, 23] or coordination of a swarm of drones [22]. The proposed method extends the experimental validation of the FIC superimposition properties to multiple redundant manipulators, allowing two arms to complete two bi-manual tasks during unknown interaction. The validation of these properties substantially simplifies the algorithmic complexity of the robot interaction controllers deployed in teleoperation and manipulation. Currently, multi-arm manipulation requires the simultaneous observation of the state of both arms through the grasp matrix, which couples the state of the two systems and increasing the algorithmic complexity [25, 26]. In contrast, the proposed architecture allows solving the same problem just by considering the simplified dynamics generated by the FIC at the end-effector.
@@ V Results
Coordination of multiple arms using synchronous via-points. |
@@
Stability often requires intricate design and tuning, and mostly by running experiments to find the stability regions. |
@@
Robust and compliant in highly variable environments without the requirement of controller gain re-tuning. |
Might require controller gain re-tuning in changing environments and various scenarios. |
@@
Might require additional energy tanks or mechanisms for ensuring stability by degrading the robot performance. |
@@
The error recorded from two cooperation experiments for drilling are the experiments in Fig. 4. They indicate that we can expect sub-millimetre accuracy when the user is holding the drill. The experiment 5.1 proves that the two arm movement can be synchronised issuing coordinated via-points without relying on trajectory planning. The results described in Fig. 5 indicate that it is possible to pick up the work-piece, carry it to the user, providing sufficient support to execute an accurate drilling, and retain stability and safety of interaction even in the event of a sudden loss of contact. The experiment 5.2 confirms that the robot can constrain the dremel inclination to the predetermined range of in each direction when drilling on a curve surface, as shown in Fig. 3. The bi-manual manipulation task trajectories (Exp. 6) are shown in Fig. 6. The arms start in the first via point and move to third to pick-up the box. The empty box is brought in to the 6 via point where the human loads the box with about of material. Regardless of the added mass the robots retain a sufficient position accuracy that enables them to carry the box back to the 3 via point, which is bounded by the FIC force saturation to a value that is . It is worth noting that the FIC is generating the maximum effort (i.e., forces and torques) selected for a task, and any position error greater than implies that the current load exceed the controller/robot capabilities. Thus, it will either require a change in the selected maximum effort in the controller, or a change of robot when the load exceed its mechanical capabilities. Experimental video is accessible through https://youtu.be/tHZ806Tjdb0.
@@ VI Discussion
The tasks show that the method can exploit its nonlinear impedance to generate higher interaction forces and accurate tracking without need of accurate modelling of the interaction dynamics. This characteristics renders this method extremely robust because it does not requires the knowledge of the environmental dynamics to stabilise its behaviour. Therefore, it presents all of the benefits that constant impedance controllers have on variable impedance and admittance controllers, but it does not suffer their low accuracy [17, 10, 1, 27]. Such accomplishment is made possible by the tunable non-linear impedance profile that enables to select the required task precision, the maximum effort and/or constant stiffness that the controller produce. Moreover, the selection of these characteristics also upper-bounds the system maximum power and maximum kinetic energy as described in [24, 22]. The nonlinear impedance allows to define low impedance funnels that can be used to guide the task performed by the user, as it has been shown in Exp. 5.2 when the user was drilling a curved work-piece. This capability might find application in training on new operators, guiding them on specific action patters, or respecting constraints that are otherwise difficult to respect (e.g. drilling with a constant angle on a curve surface).
Furthermore, it is worth noting that the interaction with the robot during the task felt very natural and all the tasks required were successful at the first attempt, and the motor synchronisation was obtained only via the mechanical interaction between the robot and the users. Finally, to better understand the unique contributions and advantages of the TeleCoop-FIC, we present a direct comparison with traditional variable impedance controllers that can be seen in Table I. TeleCoop-FIC can be considered for applications in manufacturing, space robotics, medical robotics and rehabilitation.
@@ VII Conclusion
In Conclusion, the main observed limitation is that the task accuracy and precision should be in the order of millimetre based on our set-up, but it will vary based on the robots used. Especially considering that the tuning of the controller is straight forward and can be performed by non-technical operators, simply selecting the desired task precision (), maximum effort () with respect to the physical limitation of the robot, and constant stiffness (). However, an identification of the maximum parameters is required to limit the parameter range to be within the mechanical capabilities of the robot. An identification procedure for this type of controller has been already developed, tested on multiple robots and is detailed in [17].
Future research should prioritize refining the controller’s tuning identification procedure, making it adaptable across a diverse range of robot types. Emphasis should also be placed on enhancing the system’s accuracy and precision, addressing the current millimeter-order precision limitations. Additionally, the system’s vast potential in sectors like manufacturing, space robotics, medical robotics, rehabilitation, and training beckons dedicated exploration.